Solve each system by the substitution method.
X+6y=-2
5x-20y=5

Answer by ewatrrr(24785)


Hi
Solving by the substitution method
x+6y=-2 OR x = -6y-2
5x - 20y = 5
5(-6y-2) - 20y = 5
-30y -10 - 20y = 5
-50y -10 = 5
-50y = 15
y = 15/-50 = -3/10 and
x = -2/10 = -1/5 (x = 18/10 - 20/10 )
CHECKING our Answer***
5x - 20y = 5
5*(-1/5)-20*(-3/10) = -1 + 6 = 5